Work out the following: a) 3/4 - 7/9 + 2/3

b) Which is the largest of the following numbers? 0.1, 0.02, 0.003, 0.0004, 0.00005


1
Expert's answer
2021-10-21T13:16:55-0400

a)

"\frac{3}{4}-\frac{7}{9}+\frac{2}{3}\\\frac{3}{4}-\frac{7-6}{9}\\\frac{3}{4}-\frac{1}{9}\\\frac{27-4}{36}\\=\frac{23}{36}"


b)

We change the decimal to percentage form

"0.1=\frac{1}{10}×100=10\%\\

0.02=\frac{1}{50}×100=2\%\\

0.003=\frac{3}{1000}×100=0.3\%\\

0.0004=\frac{1}{2500}×100=0.04\%\\

0.00005=\frac{1}{20000}×100=0.005\%\\"

Therefore 0.1 is the greatest because it has the highest percentage.